OpenStudy (anonymous):

find the reference angle for the given angle: 780 degrees

OpenStudy (anonymous):

please help o-o'

OpenStudy (callisto):

780 = 360 + 360 + 60 Reference angle = 60 :P

OpenStudy (anonymous):

so if the angle is bigger than 360 i subtract? o.o

OpenStudy (callisto):

For example?

OpenStudy (anonymous):

810 degrees-360=450-360=90 so would 90 be the reference angle?

OpenStudy (callisto):

Oh... 810 = 360 + 360+ 90 Reference angle= 90

OpenStudy (anonymous):

ooooh okie dokie i get this. but i had a question about negative angles too like: -199 degrees i'm confused on how to solve for it

OpenStudy (callisto):

-199 = 360-199 = 161 = 180 - 161 = 19 <- reference angle

OpenStudy (callisto):

-x = negative a angle is equal to 360 + (-x) = 360 -x

OpenStudy (anonymous):

still a little confused ^^'

OpenStudy (callisto):

Which part?

OpenStudy (anonymous):

the last part

OpenStudy (callisto):

Far to my knowledge, reference angle is angle less than /equal to 90

OpenStudy (anonymous):

ooh ok thx ^_^

OpenStudy (callisto):

and angle in quad II (that is 90<x<180) reference angle = 180-x angle in quad III (that is 180<x<270) reference angle = 180+x angle in quad IV (that is 270<x<360) reference angle = 360 - x

OpenStudy (anonymous):

ok thx i got it

OpenStudy (callisto):

Welcome :)